if there was an average intensity of 10^-6 W/M^2 measured at 10 meters from a source, what would be the sound power of that source
Answer #1

sound power = Intensity * covered area

P = I*A

= I*4pi*r^2

= 10^-6 * 4pi*10^2

= 1.25*10^-3 W

so the sound power from the source is 1.25*10^-3 W